To travel to school, Jill must walk to the tram stop. The tram trip covers four times the distance that she walks. She then need to walk 500 m to school. If the school is 10 km away from Jill's house, how far does she walk to the tram stop?
Answer by alicealc(293)

Distance she walks to the tram stop = d
Tram trip = 4d

d + 4d + 500 m = 10 km
5d + 500 m = 10000 m
5d = 10000 m - 500 m
5d = 9500 m
d = 9500/5 m = 1900 m

So, Jill walks 1900 m to the tram shop.